OpenStudy (anonymous):

(x+2)^4 + 12 = 45 I need help for solving for x's smaller and larger value :)

OpenStudy (anonymous):

\[\pm \sqrt[4]{33}-2\]

OpenStudy (anonymous):

or \[\pm33^{1/4}-2\]

hero (hero):

cinar, I did make a mistake, but you could have at least posted all your steps since you volunteered to post a solution.

OpenStudy (anonymous):

\[(x+2)^{4}+12=45\] \[(x+2)^{4}+12-12=45-12\] \[(x+2)^{4}=33\] \[((x+2)^{4})^{\frac{1}{4}}=33^{\frac{1}{4}}\] \[x+2=\pm \sqrt[4]{33}\] \[x=\pm \sqrt[4]{33}-2\]